OpenStudy (anonymous):

The current in a simple electrical circuit is inversely proportional to the resistance. If the current is 20 amperes ( an ampere is a unit for measuring current) when the resistance is 3 ohms, find the current when the resistance is 7.4 ohms.

OpenStudy (anonymous):

Oh my apologies it was on my math hw so i posted it here lol

OpenStudy (anonymous):

hmm yeah it's kinda

OpenStudy (anonymous):

3 - 1/20 7.4 - 1/x 3/x=7.4/20 x=3*20/7.4

OpenStudy (anonymous):

it says it is supposed to be a proper fraction.

OpenStudy (anonymous):

can't you convert it to proper function yourself? :)

OpenStudy (anonymous):

yes i can im just no good at it so i might get it wrong lol... and i usually just post problems i got wrong on here so i prolly wrote it wrong lol

OpenStudy (anonymous):

post what you got and i will check if it's correct :P

OpenStudy (anonymous):

so im just solving for that x up there and posting what i get just making sure haha

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok here it goes haha

OpenStudy (anonymous):

so what you got?

OpenStudy (anonymous):

i got 8.1081 repeating what is that as a fraction?

OpenStudy (anonymous):

don't use calculator :P

OpenStudy (anonymous):

im no good at making decimals into fractions i suck at mental math haha

OpenStudy (anonymous):

use paper! :P

OpenStudy (anonymous):

I meant to say i suck at all math haha

OpenStudy (anonymous):

\[7.4=\frac{74}{10}\] and you can even simplify it

OpenStudy (anonymous):

haha... wait i thought i was simplifying the number i got?

OpenStudy (anonymous):

that = 37/5

OpenStudy (anonymous):

now what haha

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

3*20=60

OpenStudy (anonymous):

\[\frac{60}{7.4}=\frac{60}{\frac{37}{5}}\]

OpenStudy (anonymous):

so now what do i do?

OpenStudy (anonymous):

do you know how to flip?

OpenStudy (anonymous):

No i do not... sorry

OpenStudy (anonymous):

you can change division by multiplication and inversing it

OpenStudy (anonymous):

\[\frac{60}{7.4}=\frac{60}{\frac{37}{5}}=60\cdot\frac{5}{37}\]

OpenStudy (anonymous):

you notice how I do it?

OpenStudy (anonymous):

so multiply that... so i got it wrong by doing it on the calculator

OpenStudy (anonymous):

no, you got same answer

OpenStudy (anonymous):

yes i do... and i did the multiplication and came up with the same answer

OpenStudy (anonymous):

\[\frac{14}{\frac{2}{7}}\] solve this yourself so you broaden your skills :P

OpenStudy (anonymous):

ok haha how do i solve it... by flipping it again or just solve as is?

OpenStudy (anonymous):

yeah by flipping

OpenStudy (anonymous):

ok so it would be 14*2/7 correct?

OpenStudy (anonymous):

I got the answer 4... is that right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

dang i tried :/ what did i do wrong?

OpenStudy (anonymous):

you didn't flip, you just changed to multiply

OpenStudy (anonymous):

didn't inverse

OpenStudy (anonymous):

Oh so it is supposed to be 14* 7/2 my bad.